Using a financial calculator please.
Today, Ava Crowder signed loan papers agreeing to borrow $4,954.85 at 9% compounded monthly. The loan payment is $143.84 a month. How many loan payments must Ava make before the loan is paid in full? A. 36 B. 39 c.) 40 D. 41 Lester Nygaard is trying to sell you an annuity that costs $184,000 today. By buying this annuity, Nygaard promises that you will receive payments of $980 per month for the next 30 years. What is the annual rate of return on this investment? A. 4.47% B. 4.93% C. 5.45% D. 5.67% s2 . Tara Knowles buys an annuity that will pay her $24,000 a year for 25 years. The payments are paid on the first day of each year. What is the value of this annuity today if the discount rate is 8.5%? A. $245,621 B. $251,409 C. $258,319 D. $266,498
Answer #1

1)

using financial calculator as

PV=-4954.85

FV=0

PMT=143.84

I/Y=9/12=0.75

Click CPT

Click N=40 Months is answer..

2)

PV=-184000

PMT=980

N=30*12=360 months

FV=0

Click CPT

Click I/Y=0.4109% per month

annual rate=0.4109%*12=4.93%

3)

press bgn button first

then

PMT=24000

N=25

I/Y=8.5

FV=0

Click CPT

Click PV=245621

the above are answers..